### v1

@@ 2.2 Singularity analysis: Fubini reduction algorithm
$(g_{1},g_{2},\ldots,g_{2L-1},g_{2L},g_{2L+1},\ldots,g_{2L+E})=(\mathcal{G}_{1},\tilde{\mathcal{G}}_{1},\ldots,\mathcal{G}_{L},\tilde{\mathcal{G}}_{L},x_{1},\ldots,x_{E})\,.$ (2.20)
@@
where $\nabla=\left[\partial_{x_{1}}\;\partial_{x_{2}}\;\ldots\;\partial_{x_{N}}\right]^{\intercal}$, and ask for what values of $\vec{s}$ the rank of $J$ changes; see, e.g., [43, Ch. 9]. In other words, we are asking for at least one null vector $\vec{\beta}=\left[\beta_{1},\beta_{2},\ldots,\beta_{k}\right]$ of $J$. (When applied to the loop-momentum representation (2.1) with $g_{i}=q_{i}^{2}-m_{i}^{2}$, $\beta$’s have the physical interpretation of Schwinger proper times). This requirement can be written as the following system of equations:
@@
where $\vec{x}\in\mathcal{U}$. We call the codimension-1 part of the solution set $\mathcal{S}(\vec{s})$ and write it as
@@
Singularity analysis of this kind applied to Feynman integrals dates back to the pioneering work of Bjorken [8], Landau [9], and Nakanishi [10] who wrote systems of polynomial equations (2.23) in various parametrizations. These are nowadays known as Landau equations and serve as a prototype for the analysis of singularities. This analysis was made rigorous through the work of Pham, Thom, Whitney, and others with the application of algebraic geometry tools to finite Feynman integrals; see [44] for a review. While mathematically rigorous, these methods tend to explode in computational complexity in anything but the simplest examples; see, e.g., [45] for a recent implementation in Macaulay2. Hence, recent resurgence of interest in this topic aims to make the singularity locus more computable.
@@
Let us briefly outline the scope of numerical tools that are currently available for cutting-edge Feynman integrals. The Julia package PLD.jl [15, 16] uses techniques from toric and polyhedral geometry to give a computable subset of the singularity locus $\mathcal{S}$. It also implements the Euler characteristic test, which is another way to probe the topology of the space $\mathcal{U}$. If for a given set of parameters $\vec{s}_{\ast}$, the absolute value of the Euler characteristic of
@@
However, the package PLD.jl is based on toric geometry and hence uses =$\mathbb{C}^{\ast}=\mathbb{C}\setminus\{0\}$ instead of $\mathbb{C}$. This is appropriate for parametric representations of Feynman integrals, but in the Baikov representation it has the effect of introducing all $\{x_{i}=0\}$ in the hypersurface union $\mathcal{U}$. This is equivalent to putting ISPs with negative exponents.
@@
Let us describe the Fubini reduction algorithm, which is a variation of [46, Sec. 4.2]. It circumvents the need to solve systems of equations such as (2.23). Instead, it proceeds by eliminating one variable $x_{i}$ at a time. We thus construct the sets of polynomials $\mathcal{L}$, $\mathcal{L}_{x_{i}}$, $\mathcal{L}_{x_{i},x_{j}}$, etc. recursively, where the subscript refers to the variables we already eliminated. At each step, we define
$[\mathcal{L}_{x_{i},\ldots,x_{j}}]_{x_{k}}\equiv\bigcup_{g\,\in\,\mathcal{L}_{x_{i},\ldots,x_{j}}}\!\!\!\!\{D_{x_{k}}(g),\,R_{x_{k}}(\infty,g)\}\;\cup\bigcup_{g,\hat{g}\,\in\,\mathcal{L}_{x_{i},\ldots,x_{j}}}\!\!\!\!\{R_{x_{k}}(g,\hat{g})\}\,.$ (2.26)
On the right-hand side, $D_{x_{k}}(g)$ refers to the discriminant of $f$ with respect to $x_{k}$, and $R_{x_{k}}(g,\hat{g})$ is the resultant of $g$ and $\hat{g}$ with respect to $x_{k}$. Writing the polynomial $g=c_{0}+c_{1}x_{k}+\ldots+c_{d}x_{k}^{d}$, the resultant $R_{x_{k}}(\infty,g)=c_{d}$ equals to the coefficient of the highest power. The set $[\mathcal{L}_{x_{i},\ldots,x_{j}}]_{x_{k}}$ might consist of reducible polynomials or repeated entries. We hence define $\mathcal{L}_{x_{i},\ldots,x_{j},x_{k}}=\text{irred}[\mathcal{L}_{x_{i},\ldots,x_{j}}]_{x_{k}}$ containing only the unique irreducible polynomials.
The three contributions in (2.26) have the following interpretations. $D_{x_{k}}(g)$ corresponds to self-pinching of $g=0$, $R_{x_{k}}(\infty,g)$ gives the pinch at infinity, while $R_{x_{k}}(g,\hat{g})$ is associated with a pinch between two curves $g=0$ and $\hat{g}=0$.
@@
We refer to the difference between $\mathcal{L}_{x_{1},x_{2},\ldots,x_{N}}$ and $\mathcal{S}$ as fictitious singularities. They are artifacts of the reduction procedure and have no physical meaning. In general, we are not concerned with catching a few fictitious singularities, since the goal is to find a superset of $\mathcal{S}$ anyway. However, it is beneficial to minimize their number.
@@
where each [$[\mathcal{L}_{\{x_{1},\ldots,x_{k-1},x_{k}\}\setminus x_{i}}]_{x_{i}}$ is defined as in (2.26). The central difference between (2.29) and (2.30) is that the former intersects $N!$ sets of proposed singularities at the very end, while the latter does so at the intermediate stages. This comes with enormous computational benefit and provides a slimmer superset of the singular locus $\mathcal{S}$, often coinciding with $\mathcal{S}$. The equation (2.30) is implemented in [Uncaptioned image] as the default solver option FastFubini.
Brown [32] and Panzer [47] have also proposed more sophisticated versions of the Fubini algorithm known as compatibility graph reduction. They are implemented in the Maple package HyperInt [14]; see [16, App. A] for a review. We found that our implementation of this approach was slower than (2.30) in practical applications, and is therefore not shipped with v1.0.0 of [Uncaptioned image] .
@@
Other approaches to determining symbol alphabets based on Landau analysis exist; for example, utilizing momentum twistor techniques or Schubert analysis; see, e.g., [48, 49, 50] for recent developments and more exhaustive lists of references. Out of the algorithms implemented as computer codes, the Mathematica package Baikovletter [51, 52, 18] uses Baikov representations to attempt constructing symbol letters. Instead of a reduction algorithm, it aims to return a list of candidate symbol letters directly. There are instances in which the resulting alphabet is not complete; see, e.g., [53, 54].
@@ 2.3 From singularities to letters
If a Feynman integral is sufficiently simple, it can admit a representation as a linear combination of multiple polylogarithms. In such cases, there is a great interest in knowing a priori the list of possible symbol letters $W_{i}$, which translate to arguments of these polylogarithms; see, e.g., [55, 56]. The collection of all independent letters is known as the symbol alphabet $\mathbb{A}$.
Before we move on, we note that even after identifying that a family of Feynman integrals includes integrals evaluating to non-polylogarithmic functions, knowledge of $\mathbb{A}$ remains valuable. This is because integrals in subsectors typically remain polylogarithmic (see [23, 57, 29] for explicit examples).
@@
where $P_{i}=P_{i}(\vec{s})$, $Q_{i}=Q_{i}(\vec{s})$, and $R_{i}=R_{i}(\vec{s})$ are some homogeneous polynomials in $\vec{s}$, such that the mass dimensions match. We assume that $Q_{i}$ is square-free, so that it does not have any factors outside the square root. This is of course just an ansatz, and one can extend it to more complicated forms if necessary. We can now apply the same analysis as before. First of all, $W_{i}$ itself has a singularity at $Q_{i}=0$. Therefore, we must have
@@
Therefore, knowing the symbol alphabet is not essential for constructing an ansatz for the differential equation (1.1) based solely on the Landau singularities $f_{j}$. However, such approach would lead to an unwieldy polynomial in the numerator. The symbol alphabet provides a practical way to reduce this ansatz into a more manageable form.
@@
There are a few different strategies to finding odd letters. To our knowledge, they were first explored in [58]. We will follow [59, Sec. 3.4], since this is the algorithm implemented in the publicly available package Effortless [36, 60]. The package assumes that a list of all candidate polynomials $Q_{i}$, $R_{i}$ in (2.32) and constants $c_{i}^{\prime}$ from (2.34) are given. It then attempts to construct all polynomials $R_{i}^{2}Q_{i}+c_{i}^{\prime}\prod_{j=1}^{K}f_{j}^{e_{ij}^{\prime}}$ that are a perfect square $P_{i}^{2}$. This can be done by scanning over all choices of $e_{ij}^{\prime}$. There is only a finite number of choices, since the set of singularities $f_{j}$ is finite and the exponents are limited because the whole expression has to have a uniform mass dimension. Repeating the same exercise for all choices of $Q_{i}$, $R_{i}$, and $c_{i}^{\prime}$ results in the proposed odd letters. Further checking for linear dependence, one obtains a candidate odd alphabet $\mathbb{A}^{\text{odd}}$ and hence also $\mathbb{A}=\mathbb{A}^{\text{even}}\cup\mathbb{A}^{\text{odd}}$.
@@
There is a handful of examples of Feynman integrals in which letters more complicated than the ansatz (2.32) exist, for example those with nested square roots [61, 62, 63]. These are currently not implemented in Effortless.
@@ 3.1 Installation
As mentioned above, [Uncaptioned image] is hosted in a GitHub repository at [3], where a compressed file with the latest version can be downloaded or cloned. The package requires a working installation of Mathematica. It was developed and tested using the Mathematica versions 13.2–14.0 and can be loaded by running the command:
@@
When an Internet connection is available, the Effortless package is automatically loaded from GitHub [60], ensuring that [Uncaptioned image] always uses the latest version. If needed, users can overwrite this path with SOFIAoptionEffortlessPath = "path/to/Effortless.m" to point to, say, a locally cloned copy. To speed up the alphabet reconstruction with Effortless, we recommend enabling FiniteFlow [64] by setting SOFIAoptionFiniteFlow = True before running Get["SOFIA.m"]. Similarly, the Julia dependencies can be activated by setting SOFIAoptionJulia = True, although this option is typically unnecessary (more on that below). A short guide for setting up an optional link between Mathematica and Julia can be found in App. B.
@@ 3.2 Summary for people with busy schedules
The code automatically assigns independent Mandelstam invariants of the form sij… $=(p_{i}+p_{j}+\ldots)^{2}$ in the cyclic basis. For example, $6$-particle Mandelstam invariants are s12, s23, s34, s45, s56, s16, s123, s234, s345 (we work in $\mathrm{D}$ dimensions). Alternatively, one can run the command:
@@
When the FindLetters -> True option is enabled, SOFIA[] first extracts the kinematic singularities and then uses the publicly available package Effortless [60] to generate all candidate letters, both even and odd. The structure of the output is then a nested list, with independent even and odd letters in the first and second entries, and square roots in the third entry, respectively:
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 2.2 Singularity analysis: Fubini reduction algorithm
$(g_{1},g_{2},\ldots,g_{2L-1},g_{2L},g_{2L+1},\ldots,g_{2L+E})=(\mathcal{G}_{1},\tilde{\mathcal{G}}_{1},\ldots,\mathcal{G}_{L},\tilde{\mathcal{G}}_{L},x_{1},\ldots,x_{E})\equiv\mathcal{L}\,.$ (2.20)
@@
where $\nabla=\left[\partial_{x_{1}}\;\partial_{x_{2}}\;\ldots\;\partial_{x_{N}}\right]^{\intercal}$, and ask for what values of $\vec{s}$ the rank of $J$ changes; see, e.g., [Cox2015, Ch. 9]. In other words, we are asking for at least one null vector $\vec{\beta}=\left[\beta_{1},\beta_{2},\ldots,\beta_{k}\right]$ of $J$. (When applied to the loop-momentum representation (2.1) with $g_{i}=q_{i}^{2}-m_{i}^{2}$, $\beta$’s have the physical interpretation of Schwinger proper times). This requirement can be written as the following system of equations:
@@
where $\vec{x}\in\mathcal{U}$, where $\mathcal{U}$ is defined in (2.19). We call the codimension-1 part of the solution set $\mathcal{S}(\vec{s})$ and write it as
@@
Singularity analysis of this kind applied to Feynman integrals dates back to the pioneering work of Bjorken [Bjorken:1959fd], Landau [Landau:1959fi], and Nakanishi [10.1143/PTP.22.128] who wrote systems of polynomial equations (2.23) in various parametrizations. These are nowadays known as Landau equations and serve as a prototype for the analysis of singularities. This analysis was made rigorous through the work of Pham, Thom, Whitney, and others with the application of algebraic geometry tools to finite Feynman integrals; see [pham2011singularities] for a review. While mathematically rigorous, these methods tend to explode in computational complexity in anything but the simplest examples; see, e.g., [Helmer:2024wax] for a recent implementation in Macaulay2. Hence, recent resurgence of interest in this topic aims to make the singularity locus more computable.
@@
Let us briefly outline the scope of numerical tools that are currently available for cutting-edge Feynman integrals. The Julia package PLD.jl [Fevola:2023kaw, Fevola:2023fzn] uses techniques from toric and polyhedral geometry to give a computable subset of the singularity locus $\mathcal{S}$. It also implements the Euler characteristic test, which is another way to probe the topology of the space $\mathcal{U}$. If for a given set of parameters $\vec{s}_{\ast}$, the absolute value of the Euler characteristic of
@@
However, the package PLD.jl is based on toric geometry and hence uses $\mathbb{C}^{\ast}=\mathbb{C}\setminus\{0\}$ instead of $\mathbb{C}$. This is appropriate for parametric representations of Feynman integrals, but in the Baikov representation it has the effect of introducing all $\{x_{i}=0\}$ in the hypersurface union $\mathcal{U}$. This is equivalent to putting ISPs with negative exponents.
@@
Let us describe the Fubini reduction algorithm, which is a variation of [Brown:2008um, Sec. 4.2]. It circumvents the need to solve systems of equations such as (2.23). Instead, it proceeds by eliminating one variable $x_{i}$ at a time. We thus construct the sets of polynomials $\mathcal{L}$, $\mathcal{L}_{x_{i}}$, $\mathcal{L}_{x_{i},x_{j}}$, etc. recursively, where the subscript refers to the variables we already eliminated. At each step, we define
[…]
Figure 2: (Left) Two-dimensional slicing of $\mathbb{C}^{N}$, with the current variable to eliminate placed along the horizontal axis and the remaining variables along the vertical axis. The curves represent the zero loci of two elements in $\mathcal{L}_{x_{i},...,x_{j}}$ we call $g=0$ and $\hat{g}=0$. Geometrically, (2.26) means that it suffices to know the special points where the roots of these polynomials (black dots) collide, indicated by the star and the white circle (detected by the discriminant) as well as the triangle (detected by the resultant). (Right) The white circle is a fictitious singularity that arises solely from the choice of coordinates. For instance, horizontal and vertical slicings lead to different fictitious singularities.
@@
We refer to the difference between $\mathcal{L}_{x_{1},x_{2},\ldots,x_{N}}$ and $\mathcal{S}$ as fictitious singularities. They are artifacts of the reduction procedure and have no physical meaning. In general, we are not concerned with catching a few fictitious singularities, since the goal is to find a superset of $\mathcal{S}$ anyway. However, it is beneficial to minimize their number. Such a fictitious point is illustrated by the circle in Fig. 2, which makes it clear that its location is potentially sensitive to the order in which the variables are eliminated (i.e., to how the slicing of $\mathbb{C}^{N}$ is performed).
@@
where each $\mathcal{L}_{\{x_{1},\ldots,x_{k-1},x_{k}\}\setminus x_{i}}]_{x_{i}}$ is defined as in (2.26). The central difference between (2.29) and (2.30) is that the former intersects $N!$ sets of proposed singularities at the very end, while the latter does so at the intermediate stages. This comes with enormous computational benefit and provides a slimmer superset of the singular locus $\mathcal{S}$, often coinciding with $\mathcal{S}$. The equation (2.30) is implemented in [Uncaptioned image] as the default solver option FastFubini.
Brown [Brown:2009ta] and Panzer [Panzer:2015ida] have also proposed more sophisticated versions of the Fubini algorithm known as compatibility graph reduction. They are implemented in the Maple package HyperInt [Panzer:2014caa]; see [Fevola:2023fzn, App. A] for a review. We found that our implementation of this approach was slower than (2.30) in practical applications, and is therefore not shipped with v1.0.0 of [Uncaptioned image] .
@@
Other approaches to determining symbol alphabets based on Landau analysis exist; for example, utilizing momentum twistor techniques or Schubert analysis; see, e.g., [Morales:2022csr, Lippstreu:2023oio, He:2024fij] for recent developments and more exhaustive lists of references. Out of the algorithms implemented as computer codes, the Mathematica package Baikovletter [Jiang:2023qnl, Jiang:2023oyq, Jiang:2024eaj] uses Baikov representations to attempt constructing symbol letters. Instead of a reduction algorithm, it aims to return a list of candidate symbol letters directly. There are instances in which the resulting alphabet is not complete; see, e.g., [Abreu:2024yit, Henn:2025xrc].
@@ 2.3 From singularities to letters
If a Feynman integral is sufficiently simple, it can admit a representation as a linear combination of multiple polylogarithms. In such cases, there is a great interest in knowing a priori the list of possible symbol letters $W_{i}$, which translate to arguments of these polylogarithms; see, e.g., [Goncharov:2010jf, Duhr:2011zq, Lee:2024kkm]. The collection of all independent letters is known as the symbol alphabet $\mathbb{A}$.
Before we move on, we note that even after identifying that a family of Feynman integrals includes integrals evaluating to non-polylogarithmic functions, knowledge of $\mathbb{A}$ remains valuable. This is because integrals in subsectors typically remain polylogarithmic (see [Bogner:2019lfa, Muller:2022gec, Giroux:2024yxu] for explicit examples).
@@
where $P_{i}=P_{i}(\vec{s})$, $Q_{i}=Q_{i}(\vec{s})$, and $R_{i}=R_{i}(\vec{s})$ are some homogeneous polynomials in $\vec{s}$, such that the mass dimensions match. We assume that $Q_{i}$ is square-free, so that it does not have any factors outside the square root. This is of course just an ansatz, and one can extend it to more complicated forms if necessary. We can now apply the same analysis as before. First of all, $W_{i}$ itself has a singularity (square-root branch point) at $Q_{i}=0$. Therefore, we must have
@@
Therefore, knowing the symbol alphabet is not essential for constructing an ansatz for the differential equation (1.1) based solely on the Landau singularities $f_{j}$. However, such an approach would lead to an unwieldy polynomial in the numerator. The symbol alphabet provides a practical way to reduce this ansatz into a more manageable form.
@@
There are a few different strategies to finding odd letters. To our knowledge, they were first explored in [Heller:2019gkq]. We will follow [Matijasic:2024too, Sec. 3.4], since this is the algorithm implemented in the publicly available package Effortless [Effortlessxxx, repoEffortless]. The package assumes that a list of all candidate polynomials $Q_{i}$, $R_{i}$ in (2.32) and constants $c_{i}^{\prime}$ from (2.34) are given. It then attempts to construct all polynomials $R_{i}^{2}Q_{i}+c_{i}^{\prime}\prod_{j=1}^{K}f_{j}^{e_{ij}^{\prime}}$ that are a perfect square $P_{i}^{2}$. This can be done by scanning over all choices of $e_{ij}^{\prime}$. There is only a finite number of choices, since the set of singularities $f_{j}$ is finite and the exponents are limited because the whole expression has to have a uniform mass dimension. Repeating the same exercise for all choices of $Q_{i}$, $R_{i}$, and $c_{i}^{\prime}$ results in the proposed odd letters. Further checking for linear dependence, one obtains a candidate odd alphabet $\mathbb{A}^{\text{odd}}$ and hence also $\mathbb{A}=\mathbb{A}^{\text{even}}\cup\mathbb{A}^{\text{odd}}$.
@@
There is a handful of examples of Feynman integrals in which letters more complicated than the ansatz (2.32) exist, for example those with nested square roots [FebresCordero:2023pww, Badger:2024fgb, Becchetti:2025oyb]. These are currently not implemented in Effortless.
@@ 3.1 Installation
As mentioned above, [Uncaptioned image] is hosted in a GitHub repository at [repo], where a compressed file with the latest version can be downloaded or cloned. The package requires a working installation of Mathematica. It was developed and tested using the Mathematica versions 13.2–14.0 and can be loaded by running the command:
@@
When an Internet connection is available, the Effortless package is automatically loaded from GitHub [repoEffortless], ensuring that [Uncaptioned image] always uses the latest version. If needed, users can overwrite this path with SOFIAoptionEffortlessPath = "path/to/Effortless.m" to point to, say, a locally cloned copy. To speed up the alphabet reconstruction with Effortless, we recommend enabling FiniteFlow [Peraro:2019svx] by setting SOFIAoptionFiniteFlow = True before running Get["SOFIA.m"]. Similarly, the Julia dependencies can be activated by setting SOFIAoptionJulia = True, although this option is typically unnecessary (more on that below). A short guide for setting up an optional link between Mathematica and Julia can be found in App. B.
@@ 3.2 Summary for people with busy schedules
The code automatically assigns independent Mandelstam invariants of the form sij… $(p_{i}+p_{j}+\ldots)^{2}$ in the cyclic basis. For example, $6$-particle Mandelstam invariants are s12, s23, s34, s45, s56, s16, s123, s234, s345 (we work in $\mathrm{D}$ dimensions). Alternatively, one can run the command:
@@
When the FindLetters -> True option is enabled, SOFIA[] first extracts the kinematic singularities and then uses the publicly available package Effortless [repoEffortless] to generate all odd candidate letters, in addition to the even letters already found by SOFIA[]. The structure of the output is then a nested list, with independent even and odd letters in the first and second entries, and square roots in the third entry, respectively:
@@
In contrast, with FindLetters -> False, the code simply returns the list of candidate singularities without generating any associated letters. As emphasized above, [Uncaptioned image] looks only for candidate singularities and a candidate symbol alphabet, meaning that it might find fictitious terms (in the sense of Sec. 2.2) that are not true singularities or true letters of a given Feynman integral. V
@@
To restrict the computation to leading singularities only, one can set IncludeSubtopologies -> False while keeping MaxCut -> True. Conversely, setting MaxCut -> False disables cuts entirely, computing both leading and subleading singularities using the top topology alone. However, this significantly increases runtime due to the much larger number of Baikov variables that must be eliminated by the polynomial solver. In this case, it is recommended to set IncludeSubtopologies -> False, as subleading singularities are already being analyzed. V
@@
Enabling Symmetries -> reflections additionally accelerates the generation of subtopologies for highly symmetric graphs, as explained in more detail below. However, because the performance gain is graph-specific, the default value is set to True instead of reflections. V
@@
The option SolverBound -> n ensures that only intermediate polynomials $\mathcal{L}_{x_{i},\ldots,x_{j}}$ in (2.26) with at most n monomials are considered in the elimination process. This option essentially controls the accuracy and speed of the computation. The higher SolverBound is, the more accurate the result becomes, but it will also take longer runtime. We recommend starting with a low value such as SolverBound -> 100. If the function returns results within a reasonable time, one should continue increasing SolverBound, as this generates a more complete list of singularities and letters if it is not already complete. The safest value is SolverBound -> Infinity. One of the few examples considered that required a high SolverBound is the “rocket diagram” from [Abreu:2024yit, Fig. 3c] covered in the notebook SOFIA_examples.nb [repo]. V